OpenStudy (anonymous):

Solve the equation: \[log_4(x)+log_7(x)=1\]

OpenStudy (anonymous):

I have no idea what to do...

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \log_{b}(x) = \frac{\log_{10}(x)}{\log_{10}(b)}\] or \[\Large \log_{b}(x) = \frac{\ln(x)}{\ln(b)}\] The second equation may be easier to work with So this means that \[\Large log_4(x)+log_7(x)=1\] is the same as \[\Large \frac{\ln(x)}{\ln(4)} + \frac{\ln(x)}{\ln(7)} = 1\]

OpenStudy (anonymous):

We did not learn the base change formula in class... so I need someone to help me out with an example first because I have no clue how to do these things... :'(

OpenStudy (mertsj):

\[\frac{\log_{10}x}{\log_{10}4}+\frac{\log_{10}x}{\log_{10}7}=\log_{10}10 \]